# The Compounded BSDE method: A fully-forward method for option pricing and optimal stopping problems in finance

### v1

@@ Abstract
Keywords: Compound BSDE method, reflected BSDE, option pricing, optimal stopping.
@@ 1 Introduction
Reflected BSDEs provide a natural framework for pricing American- and Bermudan-style options; see, for example, [23, 24, 25, 26, 19, 27] for numerical methods for solving this type of BSDE. These approaches share a similar philosophy with the least-squares Monte Carlo method [6]: one simulates the underlying dynamics forward in time and then optimizes backward. In a recent contribution [28], a backward variant of the deep BSDE method is proposed, in which the loss functional is replaced by the variance of the initial value of the $Y$ process. Numerical results suggest that this approach may be applicable to Bermudan option pricing. However, the convergence of this method for pricing early-exercise options has not been established and remains an open problem.
@@ 2.1 Review of the deep BSDE method
where $\|x\|$ denotes the Euclidean norm if $x$ is a vector or the Frobenius norm if $x$ is a matrix, $\Delta W_{t_{i}}\coloneqq W_{t_{i+1}}-W_{t_{i}}$ is the Brownian increment, $\pi$ denotes a uniform partition of $[0,T]$ with step size $h$. The quantities $\mathcal{Y}_{t_{0}}^{\pi}$ and $\mathcal{Z}_{t_{i}}^{\pi}$ are parameterized by neural networks and serve as approximations of $\mathcal{Y}_{t_{0}}$ and $\mathcal{Z}_{t_{i}}$, respectively. The objective functional (2.2a) acts as the loss function during the training of neural networks, and enforces the terminal condition $\mathcal{Y}_{T}=g(\mathcal{X}_{T})$.
@@ 3 Convergence Analysis
$\mathbb{E}\left\|g(\mathcal{X}_{T})-\mathcal{Y}_{T}^{\pi}\right\|^{2}$ $\geq(1-\epsilon^{-1}-\delta^{-1})Ch+\left[1-\epsilon+(1-\epsilon^{-1}-\delta^{-1})\left(2T^{2}K_{y}^{2}\frac{3}{1-6T^{2}K_{y}}\right)\right]\mathbb{E}\left\|\Delta\mathcal{Y}_{t_{0}}\right\|^{2}$ (3.15)
@@
$(1-\epsilon^{-1}-\delta^{-1})\left(TK_{z}^{2}+2T^{2}K_{y}^{2}\frac{3+3TK_{z}^{2}}{1-6T^{2}K_{y}^{2}}\right)+1-\delta>0,\quad 1-\epsilon+(1-\epsilon^{-1}-\delta^{-1})\left(2T^{2}K_{y}^{2}\frac{3}{1-6T^{2}K_{y}^{2}}\right)>0$ (3.16)
@@
$\hskip 170.71652pt\leq C\left(h+\sum_{j=1}^{M-1}\mathbb{E}\|g_{j}(X_{T_{j}}^{\pi},Y_{j+1,T_{j}}^{\pi})-Y_{j,T_{j}}^{\pi}\|^{2}+\mathbb{E}\|g_{M}(X_{T}^{\pi})-Y_{M,T}^{\pi}\|^{2}\right),$
@@ 5 Numerical examples
$\operatorname{Err}(Z)\coloneqq\sum_{j=1}^{M}\sum_{t_{i}\in[T_{j-1},T_{j}]}\mathbb{E}\left\|Z_{j,t_{i}}-Z_{j,t_{i}}^{\pi}\right\|^{2}h\qquad\operatorname{Total\ Err}\coloneqq\operatorname{Err}(X)+\operatorname{Err}(Y)+\operatorname{Err}(Z).$
@@
Regarding implementation, we use a fully connected neural network with two hidden layers, each with $10+d_{1}$ neurons, and $\tanh$ activations. The batch size and validation size are both $5000$, corresponding to the number of simulated sample paths used per parameter update. We train with the Adam optimizer and an exponential learning-rate decay schedule with an initial learning rate $0.01$. All computations are performed in PyTorch 2.4 on a MacBook Pro with an Apple M1 chip and 16 GB RAM.
@@ 5.1 Plain compound options with $M=2$
Table 1 presents the results for all four plain compound options. We compare the estimated price and delta at $t=0$ for $j=1$, denoted by $Y_{1,0}^{\pi}$ and $\Delta_{1,0}^{\pi}$, with the reference values $Y_{1,0}$ and $\Delta_{1,0}$ obtained from (5.3) and their analytical derivatives. We recall that the BSDE solution $(Y,Z)$ coincides with the option price and the appropriately scaled option delta via the Feynman–Kac formula, and we use this relation to recover the delta from $Z^{\pi}$. The corresponding relative mean squared errors are also reported. Overall, the method produces highly accurate estimates for both price and delta across all four contracts, with relative errors ranging from $1.04\times 10^{-3}$ down to $3.82\times 10^{-7}$.

### v2

# The Compound BSDE Method: A Fully Forward Method for Option Pricing and Optimal Stopping Problems in Finance

@@ Abstract
Keywords: Compound BSDE method, option pricing, optimal stopping, reflected BSDE.
@@ 1 Introduction
Reflected BSDEs provide a natural framework for pricing American- and Bermudan-style options; see, for example, [23, 24, 25, 26, 19, 27] for numerical methods for solving this type of BSDE. These approaches share a similar philosophy with the least-squares Monte Carlo method [6]: one simulates the underlying dynamics forward in time and then optimizes backward. In recent contributions [28, 29], a backward variant of the deep BSDE method has been studied, in which the loss functional is replaced by the variance of the $Y$ process at the initial time. Numerical results suggest that this approach may be applicable to Bermudan option pricing. However, the convergence of this method for pricing early-exercise options has not been established in [29] and remains an open problem.
@@ 2.1 Review of the deep BSDE method
where we denote $\|x\|$ the Euclidean norm if $x$ is a vector or the Frobenius norm if $x$ is a matrix, $\Delta W_{t_{i}}\coloneqq W_{t_{i+1}}-W_{t_{i}}$ is the Brownian increment, $\pi$ denotes a uniform partition of $[0,T]$ with step size $h$. The quantities $\mathcal{Y}_{t_{0}}^{\pi}$ and $\mathcal{Z}_{t_{i}}^{\pi}$ are parameterized by neural networks and serve as approximations of $\mathcal{Y}_{t_{0}}$ and $\mathcal{Z}_{t_{i}}$, respectively. The objective functional (2.2a) acts as the loss function during the training of neural networks, and enforces the terminal condition $\mathcal{Y}_{T}=g(\mathcal{X}_{T})$.
@@ 3 Convergence Analysis
$\mathbb{E}\left\|g(\mathcal{X}_{T})-\mathcal{Y}_{T}^{\pi}\right\|^{2}$ $\geq(1-\epsilon^{-1}-\delta^{-1})Ch+\left[1-\epsilon+(1-\epsilon^{-1}-\delta^{-1})\frac{6T^{2}K_{y}^{2}}{1-6T^{2}K_{y}}\right]\mathbb{E}\left\|\Delta\mathcal{Y}_{t_{0}}\right\|^{2}$ (3.15)
@@
$(1-\epsilon^{-1}-\delta^{-1})\left(TK_{z}^{2}+2T^{2}K_{y}^{2}\frac{3+3TK_{z}^{2}}{1-6T^{2}K_{y}^{2}}\right)+1-\delta>0,\quad 1-\epsilon+(1-\epsilon^{-1}-\delta^{-1})\frac{6T^{2}K_{y}^{2}}{1-6T^{2}K_{y}^{2}}>0$ (3.16)
@@
$\hskip 113.81102pt\leq C\left(h+\sum_{j=1}^{M-1}\mathbb{E}\|g_{j}(X_{T_{j}}^{\pi},Y_{j+1,T_{j}}^{\pi})-Y_{j,T_{j}}^{\pi}\|^{2}+\mathbb{E}\|g_{M}(X_{T}^{\pi})-Y_{M,T}^{\pi}\|^{2}\right),$
@@ 5 Numerical examples
$\operatorname{Err}(Z)\coloneqq\sum_{j=1}^{M}\sum_{t_{i}\in[T_{j-1},T_{j}]}\mathbb{E}\left\|Z_{j,t_{i}}-Z_{j,t_{i}}^{\pi}\right\|^{2}h,\qquad\operatorname{Total\ Err}\coloneqq\operatorname{Err}(X)+\operatorname{Err}(Y)+\operatorname{Err}(Z).$
@@
Regarding implementation, we use a fully connected neural network with two hidden layers, each with $10+d_{1}$ neurons, and $\tanh$ activations. The batch size and validation size are both $5000$, corresponding to the number of simulated sample paths used per parameter update. We train with the Adam optimizer and an exponential learning-rate decay schedule with an initial learning rate $0.01$. All computations are performed in PyTorch 2.4 on a MacBook Pro with an Apple M1 chip and 16 GB RAM. The code for the numerical experiments can be found in the GitHub repository: https://github.com/zhipeng-huang/thecompoundbsde.
@@ 5.1 Plain compound options with $M=2$
Table 1 presents the results for all four plain compound options with $N=50$. We compare the estimated price and delta at $t=0$ for $j=1$, denoted by $Y_{1,0}^{\pi}$ and $\Delta_{1,0}^{\pi}$, with the reference values $Y_{1,0}$ and $\Delta_{1,0}$ obtained from (5.3) and their analytical derivatives. We recall that the BSDE solution $(Y,Z)$ coincides with the option price and the appropriately scaled option delta via the Feynman–Kac formula, and we use this relation to recover the delta from $Z^{\pi}$. The corresponding relative mean squared errors are also reported. Overall, the method produces highly accurate estimates for both price and delta across all four contracts, with relative errors ranging from $1.04\times 10^{-3}$ down to $3.82\times 10^{-7}$.